# Lattice Protein Folding with Variational Annealing

### v1

@@ 1 Introduction & Previous Work
Protein folding is a biological process in which a linear sequence of amino acids adopts a three-dimensional structure. A correct fold or a misfold can significantly affect the biological health of a living organism [1]. As a result, an accurate understanding of how proteins fold is critical in biology and drug discovery [2]. The curse of dimensionality of the protein folding space makes it challenging to address using standard computer simulations [3]. Lattice protein folding provides a simplified yet insightful framework for studying protein folding dynamics by reducing the complexity of the search space. In the regular lattice, each cell may house an amino acid. It is also common to further simplify this folding process by reducing all 20 types of amino acids to only two types: hydrophobic and polar amino acids, also called beads. These simplifications correspond to the Hydrophobic-polar (HP) lattice protein folding model [4]. Despite these simplifications, finding the fold with the lowest energy (global minima) is NP-complete for both 2D and 3D HP lattice models [5].
@@
Machine learning tools have already addressed the question of protein folding in different settings. In the continuous folding space, AlphaFold, a machine learning-based approach has achieved remarkable performance on the prediction of protein structures compared to state-of-the-art methods [6]. In the discrete folding space, machine learning approaches have also addressed the HP protein folding model in 2D as a combinatorial optimization problem. In recent years, there has been increasing interest in using machine learning techniques to solve combinatorial optimization problems [7, 8, 9, 10, 11, 12, 13, 14, 15, 16]. In particular, for the 2D lattice protein folding problem, this study [17] developed a model called FoldingZero, which combines deep reinforcement learning (RL) with a two-head deep convolutional neural network (HPNet) and a modified tree search algorithm. This work investigated chain sizes up to $85$ with successful runs matching the optimal energy for sizes up to $20$. Another RL study [18] also investigated HP chains up to $36$ beads using various RL methods such as policy and value iteration, Monte Carlo Tree Search, and AlphaGo Zero with pretraining. According to their results, the adoption of the AlphaGo Zero algorithm exhibits superior performance in comparison to the other methods. Finally, by far the strongest RL work [19] obtains optimal folds for HP chains up to $50$ beads. They largely attribute the success of their method to the incorporation of Long Short Term Memory (LSTM) architectures in their procedure, which capture long-range interactions in the folding process.
@@
Our work demonstrates the ability of dilated recurrent neural networks (RNNs) [20] supplemented with temperature annealing  [12, 13, 14, 15] to solve instances of the 2D HP lattice folding model. In all the previous studies, sampling invalid folds has been discouraged by introducing an energy penalty. In our work, we use masking to sample valid folds from RNNs autoregressively to enhance convergence and training stability. We also introduce a novel scheme by introducing a free energy upper bound that stabilizes and enhances RNNs training on the 2D HP model, while preserving their ability to generate folds in the valid folding space autoregressively.
@@ 2.1 The HP Model
A fully folded protein chain can be conveniently represented on the 2D Cartesian plane. Let $\Gamma=(\gamma_{0},\dots,\gamma_{N})\in\{0,1\}^{N+1}$ represent an HP chain or sequence having $N+1$ beads. Here, 00 and $1$ denote the ‘H’ and ‘P’ beads respectively. For a complete fold of $\Gamma$, let the Cartesian coordinates of each bead in $\Gamma$ be $(x_{0},y_{0}),\dots,(x_{N},y_{N})$ respectively where $\forall i;x_{i},y_{i}\in\mathbb{Z}$. A fold requires every pair of consecutive beads to be a unit distance away from each other, either on the $x$-axis or on the $y-$axis, but not both. In other words, the following condition
@@
must also hold true for a valid fold. To map a protein sequence $\Gamma$ to some (valid or invalid) fold, we define a sequence of moves to be the solution $\bm{d}\in\{0,1,2,3\}^{N}$. A move $d_{i}\in\bm{d}$ dictates the coordinates of bead $\gamma_{i}$ in a fold given the coordinates of the previous bead $\gamma_{i-1}$ which are $(x_{i-1},y_{i-1})$. Specifically, for $1\leq i\leq N$, the next bead position is given as follows:
@@
The initial position $(x_{0},y_{0})$ can be set to any reference coordinates, such as $(0,0)$, which we use in this work. Semantically, our chosen convention is such that moves $0,1,2,$ and $3$ corresponds to placing the current bead $\gamma_{i}$ to the ‘left of’, ‘right of’, ‘above’, and ‘below’ the previous bead $\gamma_{i-1}$ respectively on the Cartesian plane.
@@
Here, the energy function checks all four neighboring coordinates of $\gamma_{i}$ in the fold. An energy contribution of $-1$ is added in proportion to the number of neighboring ‘H’ beads through the term proportional to $\widehat{M_{i}}$. However, for all beads, — except for the first and the last — two of the four neighboring beads $\gamma_{i-1}$ and $\gamma_{i+1}$ cannot contribute to the energy as they are consecutive to $\gamma_{i}$ in the protein sequence. Therefore, we subtract these contributions by adding the term proportional to $A_{i}$. For the boundary cases of the first and the last bead, we use $\gamma_{-1}=\gamma_{N+1}=1$. Note that the double counting of ‘H-H’ adjacent pairs is taken into account by dividing by a factor of 2. Lastly, if a fold $\boldsymbol{d}$ is invalid, its energy is set to 00 by default.
@@ 2.2 Variational Annealing
In Eq. (6), the term $T\log(P_{\boldsymbol{\theta}}(\boldsymbol{d}^{(i)}))$ can be seen as an entropy regularization term weighted by temperature $T$ [12, 24, 13, 15, 16]. We use this entropy regularization to mitigate the effects of local minima in the optimization landscape [12] and also to avoid mode collapse [24]. $T$ is annealed or cooled from a starting temperature $T_{0}$ to a final temperature 00 with the possibility of varying curvatures in its descent depending on the annealing schedule – ranging from a steady, linear decay to a faster, nonlinear decay that follows the curvature of the inverse function for example. Selecting an annealing schedule is a design choice that dictates how fast $T$ decays during the different stages of the annealing process. With these schedules, entropy regularization can be seen as a mechanism that encourages exploration in the folds landscape at high temperatures before exploitation by targeting the low energy folds near zero temperature.
@@ 2.3 Probabilistic Model
To model the probability distribution from which folds are sampled from $P_{\boldsymbol{\theta}}(\boldsymbol{d})$, we use a Dilated RNN architecture [20]. The motivation behind using an RNN architecture is to enable autoregressive sampling, which is a form of perfect sampling that mitigates the challenges of Markov Chain sampling schemes of other neural network architectures [25, 26, 27]. Furthermore, unlike the vanilla RNN model, Dilated RNNs have long recurrent skip connections that allow for the direct propagation of hidden state information from earlier inputs $\bm{x}_{i}$ to be utilized in later stages of the folding process. This property is particularly useful in the context of folding as we may want to put an ‘H’ bead $\gamma_{i}$ adjacent to an ‘H’ bead $\gamma_{j}$ where $i-j$ is large. These long-term dependencies benefit from the introduced dilated recurrent connections [20].
@@
Here, $\boldsymbol{x}_{n-1}$ is the input (to the first layer of the Dilated RNN stack) that is a concatenation of the one-hot encoding of the protein bead for which we want to sample a fold for $\boldsymbol{q}_{n}$ and the one-hot encoding of the previously sampled output fold $\boldsymbol{d}_{n-1}$. More concretely, $\boldsymbol{q}_{n}$ is a one-hot encoding vector of $\{0,1\}$ where 00 represents ‘H’ and $1$ represents ‘P’, and $\boldsymbol{d}_{n}$ is the one-hot encoding vector of integer $d_{n}\in\{0,1,2,3\}$. Using these two vectors, we can construct the input $\boldsymbol{x}_{n-1}=[\boldsymbol{q}_{n}\frown\boldsymbol{d}_{n-1}]$ where $\frown$ is the concatenation operation. Note that the initializations of the hidden state are defined as $\boldsymbol{h}_{n}^{0}=\boldsymbol{x}_{n-1}$ and $\boldsymbol{h}_{0}^{l}=[\boldsymbol{q}_{0}\frown\boldsymbol{0}]$.
@@
Finally, we would like to highlight that the computational complexity of our model for sampling a configuration or for a forward pass is $\mathcal{O}(N\log(N))$, which makes our RNN model more scalable compared to traditional Transformers [28].
@@ 2.4 Masking and Upper Bound Optimization
such that $P^{\perp}_{\boldsymbol{\theta}}$ is the masked RNN probability with a stop gradient assignment $\perp$ [12, 31]. Minimizing $\mathcal{L}(T)$ corresponds to the REINFORCE method [32] with a vanilla policy gradient rule [33, 34] supplemented with an entropy term [12, 13, 14, 15, 24, 35]. To train our Dilated RNNs, we use the following cost function:
@@
The first inequality follows from the fact that $E(\bm{d})\leq 0$ for all possible folds $\bm{d}\in\{0,1,2,3\}^{N}$. Training using the upper bound loss function ~$\tilde{\mathcal{L}}(T)$ follows a similar spirit to training variational autoencoders [36] and diffusion models [37, 38, 39] with upper bound loss functions. The gradient of the free energy upper bound is given as:
@@
Note that we used the notation $\overline{O(\boldsymbol{d})}\equiv O(\boldsymbol{d})-\langle O\rangle$, where subtracting the average of the energies and log probabilities was shown to reduce the variance in the gradients as a control variate method [33, 30, 40].
@@ 3 Experiments & Results
An overview of the training scheme of our RNN is presented in Algorithm 1 and also illustrated in Fig. 3. In our method, we use a linear schedule for the annealing process which is characterized by Line 3 in the pseudocode. Further hyperparameter details are provided in Appendix A.
@@
1:Required: $T_{0},N_{\text{anneal}},N_{\text{train}},M$
@@
5:         Sample $M$ independent folds from the Dilated RNN
@@
20merA HPHPPHHPHPPHPHHPPHPH 20 $-9$
20merB HHHPPHPHPHPPHPHPHPPH 20 $-10$
24mer HHPPHPPHPPHPPHPPHPPHPPHH 24 $-9$
25mer PPHPPHHPPPPHHPPPPHHPPPPHH 25 $-8$
36mer PPPHHPPHHPPPPPHHHHHHHPPHHPPPPHHPPHPP 36 $-14$
48mer PPHPPHHPPHHPPPPPHHHHHHHHHHPPPPPPHHPP- 48 $-23$
@@ 3.3 Benchmarks
To compare our results with the machine learning literature, we report in Tab. 1 a list of HP sequences with lengths from $20$ to $60$ and their optimal energies that are often used to benchmark HP folding algorithms [19]. We use these HP sequences to benchmark our algorithm and compare with recent machine learning methods on the criteria of the lowest energy fold found. The results are presented in Tab. 2. Our findings demonstrate that our variational annealing method with dilated RNNs can find the optimal fold for all the sequences up to $60$ beads. We also highlight that our method performs better than previous studies that take inspiration from the AlphaGo Zero algorithm [17, 18]. Our method is also competitive with the results reported by Ref. [19], where we only use a simple policy gradient rule compared to the advanced policy gradient rules used in Ref. [19]. Additionally, we note that we only use $d_{h}=50$ as the size of the hidden state in our dilated RNN compared to the LSTM used in Ref. [19] with $d_{h}\in\{256,512\}$. This observation highlights the computational efficiency of our method. Most importantly, we identify the optimal fold for the 60-bead configurations, representing a notable improvement over the machine learning optimization techniques reported in this paper (see Tab. 2).
Finally, we highlight that our RNN model finds diverse solutions at the end of annealing. In particular, Fig. 5 shows the evolution of the Shannon entropy of the RNN as a function of the number of annealing steps $N_{\rm step}$. By the end of annealing, we find non-zero final entropy values of $2.68,2.38,2.57$ for the 20merA, 25mer, and 50mer protein chains respectively. This observation highlights the ability of our RNN model to find a diverse set of solutions. We also note the entropy maxima around $T=0.096,0.121,0.394$ for the three protein chains respectively. These peaks could be related to the onset of a phase transition from a disordered phase at high temperatures (with large entropy) to another phase at low temperatures (with low entropy), which contains the low-energy solutions of the HP lattice model. We envision that slowing down the annealing schedule near the entropy peak could improve the adiabaticity of our annealing protocol, and could lead to a better heuristic solver based on our annealing framework.
Figure 5: Figures of the Shannon entropies $S_{\boldsymbol{\theta}}$ during training at every annealing stage spanning $N_{\text{annealing}}=10000$ steps on the sequences (a) 20merA, (b) 25mer, and (c) 50mer.

### v2

@@ 1 Introduction & Previous Work
Protein folding is a biological process in which a linear sequence of amino acids adopts a three-dimensional structure. A correct fold or a misfold can significantly affect the biological health of a living organism [1]. As a result, an accurate understanding of how proteins fold is critical in biology and drug discovery [2]. The curse of dimensionality of the protein folding space makes it challenging to address using standard computer simulations [3]. Lattice protein folding provides a simplified yet insightful framework for studying protein folding dynamics by reducing the complexity of the search space. In the regular lattice, each cell may house an amino acid. It is also common to further simplify this folding process by reducing all 20 types of amino acids to only two types: hydrophobic and polar amino acids, also called beads. These simplifications correspond to the Hydrophobic-polar (HP) lattice protein folding model [4], which is a model that helps elucidate fundamental principles of protein folding, such as the role of hydrophobic interactions and the relationship between amino acid sequences and native structures [5]. Despite these simplifications, finding the fold with the lowest energy (global minima) is NP-complete for both 2D and 3D HP lattice models [5].
@@
Machine learning tools have already addressed the question of protein folding in different settings. In the continuous folding space, AlphaFold, a machine learning-based approach has achieved remarkable performance on the prediction of protein structures compared to state-of-the-art methods [6]. In the discrete folding space, machine learning approaches have also addressed the HP protein folding model in 2D as a combinatorial optimization problem. In recent years, there has been increasing interest in using machine learning techniques to solve combinatorial optimization problems [7]. These techniques span different paradigms of learning such as reinforcement learning [8, 9] and unsupervised learning [10, 11]. They also incorporate a wide variety of neural network architectures such as recurrent neural networks [12, 10], graph neural networks [13, 14], and diffusion models [15]. In particular, for the 2D lattice protein folding problem, this study [16] developed a model called FoldingZero, which combines deep reinforcement learning (RL) with a two-head deep convolutional neural network (HPNet) and a modified tree search algorithm. This work investigated chain sizes up to 858585 with successful runs matching the optimal energy for sizes up to 202020. Another RL study [17] also investigated HP chains up to 363636 beads using various RL methods such as policy and value iteration, Monte Carlo Tree Search, and AlphaGo Zero with pretraining. According to their results, the adoption of the AlphaGo Zero algorithm exhibits superior performance in comparison to the other methods. Finally, by far the strongest RL work [18] obtains optimal folds for HP chains up to 505050 beads. They largely attribute the success of their method to the incorporation of Long Short Term Memory (LSTM) architectures in their procedure, which capture long-range interactions in the folding process.
@@
Our work demonstrates the ability of dilated recurrent neural networks (RNNs) [19] supplemented with temperature annealing  [10, 14] to solve instances of the 2D HP lattice folding model. In all the previous studies, sampling invalid folds has been discouraged by introducing an energy penalty. In our work, we use masking to sample valid folds from RNNs autoregressively to enhance convergence and training stability. We also introduce a novel scheme by introducing a free energy upper bound that stabilizes and enhances RNNs training on the 2D HP model, while preserving their ability to generate folds in the valid folding space autoregressively.
@@ 2.1 The HP Model
A fully folded protein chain can be conveniently represented on the 2D Cartesian plane. Let $\Gamma=(\gamma_{0},\dots,\gamma_{N})\in\{0,1\}^{N+1}$ represent an HP chain or sequence having $N+1$ beads. Here, 0 and 111 denote the ‘H’ and ‘P’ beads respectively. For a complete fold of $\Gamma$, let the Cartesian coordinates of each bead in $\Gamma$ be $(x_{0},y_{0}),\dots,(x_{N},y_{N})$ respectively where $\forall i;x_{i},y_{i}\in\mathbb{Z}$. A fold requires every pair of consecutive beads to be a unit distance away from each other, either on the $x$-axis or on the $y$-axis, but not both. In other words, the following condition
@@
must also hold true for a valid fold. To map a protein sequence $\Gamma$ to some (valid or invalid) fold, we define a sequence of moves or actions to be the solution $\bm{d}\in\{0,1,2,3\}^{N}$. A move $d_{i}\in\bm{d}$ dictates the coordinates of bead $\gamma_{i}$ in a fold given the coordinates of the previous bead $\gamma_{i-1}$ which are $(x_{i-1},y_{i-1})$. Specifically, for $1\leq i\leq N$, the next bead position is given as follows:
@@
The initial position $(x_{0},y_{0})$ can be set to any reference coordinates, such as (0,0)(0,0)( 0 , 0 ), which we use in this work. Semantically, our chosen convention is such that moves 0,1,2,0,1,2,0 , 1 , 2 , and 333 corresponds to placing the current bead $\gamma_{i}$ to the ‘left of’, ‘right of’, ‘above’, and ‘below’ the previous bead $\gamma_{i-1}$ respectively on the Cartesian plane.
@@
Here, the energy function checks all four neighboring coordinates of $\gamma_{i}$ in the fold. An energy contribution of 1-1- 1 is added in proportion to the number of neighboring ‘H’ beads through the term proportional to $\widehat{M_{i}}$. However, for all beads, — except for the first and the last — two of the four neighboring beads $\gamma_{i-1}$ and $\gamma_{i+1}$ cannot contribute to the energy as they are consecutive to $\gamma_{i}$ in the protein sequence. Therefore, we subtract these contributions by adding the term proportional to $A_{i}$. For the boundary cases of the first and the last bead, we use $\gamma_{-1}=\gamma_{N+1}=1$. Note that the double counting of ‘H-H’ adjacent pairs is taken into account by dividing by a factor of 2. Lastly, if a fold $\boldsymbol{d}$ is invalid, its energy is set to 0 by default.
@@
Figure 2: The sequence ‘PHHPHHPH’ is being folded one RNN-generated move or action $d$ at a time, for a total of $N=7$ steps. We see how each action positions the next bead relative to the previous bead in the partial fold. Note the different colors on the RNN blocks to denote position-wise dedicated parameters instead of the same set of shared parameters for all RNN blocks as specified in Sec. 2.3.
@@ 2.2 Variational Annealing
In Eq. (6), the term $T\log(P_{\boldsymbol{\theta}}(\boldsymbol{d}^{(i)}))$ can be seen as an entropy regularization term weighted by temperature $T$. We use this entropy regularization to mitigate the effects of local minima in the optimization landscape [10] and also to avoid mode collapse [22]. $T$ is annealed or cooled from a starting temperature $T_{0}$ to a final temperature 0 with the possibility of varying curvatures in its descent depending on the annealing schedule – ranging from a steady, linear decay to a faster, nonlinear decay that follows the curvature of the inverse function for example. Selecting an annealing schedule is a design choice that dictates how fast $T$ decays during the different stages of the annealing process. With these schedules, entropy regularization can be seen as a mechanism that encourages exploration in the folds landscape at high temperatures before exploitation by targeting the low energy folds near zero temperature.
@@ 2.3 Probabilistic Model
To model the probability distribution from which folds are sampled from $P_{\boldsymbol{\theta}}(\boldsymbol{d})$, we use a Dilated RNN architecture [19]. The motivation behind using an RNN architecture is to enable autoregressive sampling, which is a form of perfect sampling that mitigates the challenges of Markov Chain sampling schemes of other neural network architectures [23, 24]. Furthermore, unlike the vanilla RNN model, Dilated RNNs have long recurrent skip connections that allow for the direct propagation of hidden state information from earlier inputs $\bm{x}_{i}$ to be utilized in later stages of the folding process. This property is particularly useful in the context of folding as we may want to put an ‘H’ bead $\gamma_{i}$ adjacent to an ‘H’ bead $\gamma_{j}$ where $i-j$ is large. These long-term dependencies benefit from the introduced dilated recurrent connections [19]. Fig 2 illustrates the step-by-step procedure for generating a complete fold from the starting bead by the Dilated RNN.
@@
[…]
@@
Finally, we would like to highlight that the computational complexity of our model for sampling a configuration or for a forward pass is $\mathcal{O}(N\log(N))$, which makes our RNN model more scalable compared to traditional Transformers [25]. However, at an increased cost of complexity, transformers may better capture long-range interactions.
@@ 2.4 Masking and Upper Bound Optimization
such that $P^{\perp}_{\boldsymbol{\theta}}$ is the masked RNN probability with a stop gradient assignment $\perp$ [10, 28]. Minimizing $\mathcal{L}(T)$ corresponds to the REINFORCE method [29] with a vanilla policy gradient rule [30, 31] supplemented with an entropy term [22]. To train our Dilated RNNs, we use the following cost function:
@@
The first inequality follows from the fact that $E(\bm{d})\leq 0$ for all possible folds $\bm{d}\in\{0,1,2,3\}^{N}$. Training using the upper bound loss function ~$\tilde{\mathcal{L}}(T)$ follows a similar spirit to training variational autoencoders [32] and diffusion models [33, 34] with upper bound loss functions. The gradient of the free energy upper bound is given as:
@@
Note that we used the notation $\overline{O(\boldsymbol{d})}\equiv O(\boldsymbol{d})-\langle O\rangle$, where subtracting the average of the energies and log probabilities was shown to reduce the variance in the gradients as a control variate method [27, 30].
@@ 3 Experiments & Results
An overview of the training scheme of our RNN is presented in Algorithm 1 and also illustrated in Fig. 4. In our method, we use a linear schedule for the annealing process which is characterized by Line 3 in the pseudocode. Further hyperparameter details are provided in Appendix A, Tab. 3.
@@
1:Required: $\Gamma,T_{0},N_{\text{anneal}},N_{\text{train}},M$
@@
5:   Sample $M$ independent folds of protein $\Gamma$ from the RNN.
@@
20merA HPHPPHHPHPPHPHHPPHPH 20 9-9- 9
20merB HHHPPHPHPHPPHPHPHPPH 20 10-10- 10
24mer HHPPHPPHPPHPPHPPHPPHPPHH 24 9-9- 9
25mer PPHPPHHPPPPHHPPPPHHPPPPHH 25 8-8- 8
36mer PPPHHPPHHPPPPPHHHHHHHPPHHPPPPHHPPHPP 36 14-14- 14
48mer PPHPPHHPPHHPPPPPHHHHHHHHHHPPPPPPHHPP- 48 23-23- 23
@@
Figure 6: Some representative, optimal folds found by the VNA method for HP sequence (a) 20merA, (b) 25mer, and (c) 50mer. The solid and hollow circles represent ‘H’ and ‘P’ beads respectively. The arrow indicates the start of the fold.
@@ 3.2 Masking Ablation
Figure 7: Conditional probability distribution over the four moves for the last step $d_{19}$ of the 20merA sequence at different stages of the training process. Notice that the conditional probability of the invalid ‘left’ move is always 0 and by the end of training, all the probability mass is on the optimal ‘up’ move.
@@ 3.3 Benchmarks
To compare our results with the machine learning literature, we present in Tab. 1 a list of HP sequences with lengths from 202020 to 606060 and their optimal energies that are often used to benchmark HP folding algorithms [18]. We use these HP sequences to benchmark our algorithm and compare with recent machine learning methods on the criteria of the lowest energy fold found. The results are reported in Tab. 2 and some optimal folds that our method finds are shown in Fig. 6. Our findings demonstrate that our variational annealing method with dilated RNNs can find the optimal fold for all the sequences up to 606060 beads. We also highlight that our method performs better than previous studies that take inspiration from the AlphaGo Zero algorithm [16, 17]. Additionally, we note that we only use $d_{h}=50$ as the size of the hidden state in our dilated RNN compared to the LSTM used in Ref. [18] with $d_{h}\in\{256,512\}$. This observation highlights the computational efficiency of our method. Most importantly, we identify the optimal fold for the 60-bead configurations, representing a notable improvement over the machine learning optimization techniques reported in this paper (see Tab. 2). We performed thorough experiments on chains up to a reasonable size of 60 residues, but note that experiments on larger chains can also be carried out, albeit at higher computational costs. Furthermore, we report an example of the evolution of the conditional probability of the actions made by the RNN in Fig. 7 for the 20merA sequence. Here the confidence in making an action that increases the number of H-H interactions is increasing with more training steps, indicating the RNN is indeed learning to maximize the number of H-H connections. The low-energy states of the HP model which give rise to the structures of the folded proteins (see Fig. 6) consist of a core of H beads that exhibit high compactness. This confirms known results regarding the folded states of HP sequences [5]. Consequently, this suggests that the variational annealing method could help describe the principles underlying lattice protein folding.
Finally, we highlight that our RNN model finds diverse solutions at the end of annealing. In particular, Fig. 8 shows the evolution of the Shannon entropy of the RNN as a function of the number of annealing steps $N_{\rm step}$. By the end of annealing, we find non-zero final entropy values of 2.69,1.33,0.562.69,1.33,0.562.69 , 1.33 , 0.56 for the 20merA, 25mer, and 36mer protein chains respectively. This observation highlights the ability of our RNN model to find a diverse set of solutions.
Figure 8: Figures of the Shannon entropies $S_{\boldsymbol{\theta}}$ during training at every annealing stage spanning $N_{\text{annealing}}=10000$ steps on the sequences (a) 20merA, (b) 25mer, and (c) 36mer.